OpenStudy (anonymous):

Please help with this sigma notation question? What is the series in sigma notation? -5 + (-1) + 3 + 7 + 11?

OpenStudy (anonymous):

What's the common difference between the numbers?

OpenStudy (anonymous):

There adding 4 each time?

OpenStudy (anonymous):

Would the answer be 5 Z (4i - 9) i = 1

OpenStudy (anonymous):

Yes. You could also start at i=0 and say 4 Z (4i-5) i=0 if you wanted to, but yours works just as well. Good job!

OpenStudy (anonymous):

Thanks! Can u help me with one more to?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Which represents the series in sigma notation? 1 + 1/2 + 1/4 + 1/8 + 1/16 + 1/32 +1/64?

OpenStudy (anonymous):

The common difference would be that there all being times by 2 to get the next number

OpenStudy (anonymous):

Would the answer be 7 Z 1/2^( i + 1) i = 1

OpenStudy (anonymous):

You've got the right idea, but your exponent is wrong. When you plug in i=1, you get the answer 1/4 instead of 1.

OpenStudy (anonymous):

None of my answers have 1/4 in them. There all 1/2 or 2

OpenStudy (anonymous):